OpenStudy (anonymous):

what is the slope and x and y intercept of x-8y=3

jigglypuff314 (jigglypuff314):

Hello and Welcome to Openstudy! :) you can solve for y (get the y alone) so that it'll look like y = mx + b where then m=slope and b=y-intercept x - 8y = 3 to get y alone, minus x from both sides -8y = -x + 3 divide both sides by -8 y = 1/8 x - 3/8 ^m ^b so m=slope= and b= y-intercept =

OpenStudy (anonymous):

ok so how would I do that when it is y=7

jigglypuff314 (jigglypuff314):

hmm? what do you mean?